what are the zeros of the equation -x^3+5x^2-17x+13
1
Expert's answer
2014-07-01T06:47:33-0400
We can rewrite the equation in the next form: (x-1)(x^2-4x+13)=0. ----> x-1=0 or x^2-4x+13=0.

x^2-4x+13=(x-2)^2+9. So, the zeros of the equation x^2-4x+13: 2-3i, 2+3i.

Thus, we have that 1, 2-3i and 2+3i are zeros of the equation -x^3+5x^2-17x+13=0.